a 900kg roller coaster car starts from rest at point a(on hill) of a height,H,rolls down the track, goes around the ( points B(bottom loop) and C (top loop). Find the velocity at point C ,if the work done to move the roller coaster from point B to point C is 264870J.Given the dimensions are : H=80m, r =15m,
r(radius of loop)
1
Expert's answer
2020-07-22T10:33:38-0400

"mgH=mgh+mv^2\/2"


"mgh=W=246870J"


"v=\sqrt{2(mgH-mgh)/m}=\sqrt{2(900\cdot 9.81\cdot80-246870)/900}=31.95m/s"